Solve the following:
Find the slope of the line passing through the points (3, 4) and (3, –4).

Found 2 solutions by checkley75, madmaths:
Answer by checkley75(3666)

SLOPE=(Y2-Y1)/(X2-X1)
SLOPE=(-4-4)/(3-3)
SLOPE=-16/0 MATHEMATICALLY THIS WOULD BE INFINITY BUT INSTEAD IT IS A VERTICLE LINE PASSING THROUGH THE POINT X=3.
TO PROVE THIS FACT PLOT THE TWO POINTS & DRAW A STRAIGHT LINE THROUGH THEM.

Answer by madmaths(2)

let x1=3 and y1=4
and x2=3 and y2=-4
now slope=(y2-y1)/(x2-x1)
= (-4-4)/(3-3)
= -8/0
=infinity
if the slope is infinity it means line is parallel to y axis.